OpenStudy (anonymous):

Find the integral. Possibly with u-substitution. 2/3(Indefinite Integral of sqrt(4-3/5x)dx

OpenStudy (anonymous):

\[2/3\int\limits_{}^{}\sqrt{4-3x/5}dx\]

hartnn (hartnn):

is it \(\Large 4- \dfrac{3x}{5}\) if yes, then you can put u = that

hartnn (hartnn):

and find du

OpenStudy (anonymous):

u = 4-3x/5 du = -15 / 25 = -3/5 dx dx = 5du/-3 That's what I got.

hartnn (hartnn):

thats correct so where are you stuck now ?

hartnn (hartnn):

(2/3) (-3/5) integral sqrt u du that should be easy